OpenStudy (newarf):

if m^9 / m^6 = 8 what is the value of m ^5 / m^2

OpenStudy (peachpi):

subtract the exponents \[\frac{ m^9 }{ m^6 }=m^{9-6}=m^3\] Also, \[\frac{ m^5 }{ m^2 }=m^{5-2}=m^3\]